OpenStudy (anonymous):

When Jason started the job he had 213 1/2 pounds of nails. When he finished he had 167 3/4 pounds of nails. How many pounds of nails did he use on the job?

OpenStudy (anonymous):

HELPPPPPPPPPPPPPP!

hero (hero):

Hint: \[\space\space\space\space\space213 \frac{1}{2} \\- 167 \frac{3}{4} \\\text{_________} \]

hero (hero):

Subtract

OpenStudy (anonymous):

OKAY WAIT

OpenStudy (anonymous):

right. Now to subtract the fractions you need both denominators to be the same

hero (hero):

Yes, so find the equivalent fraction for 1/2 and replace it.

OpenStudy (anonymous):

CAN YOU GIVE THE ANSWER BECAUSE IM KIND IN A RUSH

OpenStudy (anonymous):

I HAVE TO GO TO SCHOOL PLEASE HERO!

hero (hero):

1/2 = 2/4

hero (hero):

\[\space\space\space\space\space213 \frac{2}{4} \\- 167 \frac{3}{4} \\\text{_________} \]

hero (hero):

Now all you have to do is use barrowing methods to figure it out

OpenStudy (anonymous):

45.75

hero (hero):

How did you get that? Using a calculator?

OpenStudy (anonymous):

YEAH

hero (hero):

:/

OpenStudy (anonymous):

WHAT IM IN A RUSH

OpenStudy (anonymous):